OpenStudy (anonymous):

Solve the following and Write Answer in Interval Notation. -2(x+1) -3 ≤ -12 and 3x-5 <13

OpenStudy (anonymous):

On solving the first equation you'll get \[x \ge 7/2 \] The intreval is : \[[3.5, \infty )\] On solving the second equation you'll get \[x < 6\] The intreval is : \[( -\infty , 6 )\]

OpenStudy (anonymous):

how would i put both of the answers in interval notation so that it is together?

OpenStudy (anonymous):

Oh in that case, you must find the intersection of the two intervals. The answer is [ 3.5 , 6 ) .